OpenStudy (anonymous):

what is the third term expansion of (3x-2)^5

OpenStudy (anonymous):

(3x-2)^5=(3x)^5+(-2)(3x)^4+(-2)^2*(3x)^3+(-2)^3*(3x)^2+(-2)^4*(3x)+(-2)^5 OR (3x-2)^5=-32+240 x-720 x^2+1080 x^3-810 x^4+243 x^5

OpenStudy (anonymous):

Thank you so much! :D

OpenStudy (anonymous):

you are welcome

OpenStudy (anonymous):

are you still on

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i keep getting one